Gdeinward:

What volume is equivalent to 12.0 m3? 1.2 x 10^‒6 km3 1.2 x 10^‒3 km3 1.2 x 10^4 mm3 1.2 x 10^10 mm3

Gdeinward:

ALL CREW ON DECK!!!!! I am having some trouble with this. I know the answer is C. Ive used calculators, followed exactly as the lesson says, and tried to solve it myself. However, the answer is marked as incorrect when I submitted the quiz. I am freaking out, I feel like the guy from The Raven over here.

Gdeinward:

I know im probably missing something, but I really feel like im right on this one.

Gdeinward:

@Vocaloid

Gdeinward:

@Ultrilliam

Gdeinward:

@TheSmartOne

Narad:

\[1m= 10^{3} mm\] \[1m^{3}=10^{9} mm ^{3}\] Therefore, \[12 m ^{3}= ????? mm ^{3}\]

Gdeinward:

So the answer is C? @dude

dude:

Whoops, misread Narad explained the proper way, I overlooked the fact that its cubed Its not C Youre trying to find what is equal to 12,000,000,000mm\(^3\)

TheSmartOne:

Agreed, it's not C. You have to be careful since your units are m^3 not m. 1 m = 1000 mm = 10^3 mm that's the conversion between m and mm But the conversion will be different going in between m^3 and mm^3 \(1 m^3 = (1000)^3 mm^3 = (10^3)^3 m^3 = 10^9 m^3\)

Gdeinward:

Thanks guys. I really appreciate it. I did notice how it was cubed, and I tried solving it, but got the same answer.